Use conditional probability to find each of the following.
(a) What is the probability of randomly selecting a club from a deck of cards given that the card drawn in a black card?






(b) What is the probability of randomly selecting a queen from a deck of cards given that the card drawn is a face card?


Answer by stanbon(75887)

(a) What is the probability of randomly selecting a club from a deck of cards given that the card drawn in a black card?
----
P(club | black) = P(club AND black)/P(black) = (13/52)/(26/52) = 1/2
------------------------
(b) What is the probability of randomly selecting a queen from a deck of cards given that the card drawn is a face card?
P(queen | face) = P(queen AND face)/P(face) = (4/52)/(12/52) = 1/3
=====================================================================
Cheers,
Stan H.